OpenStudy (cutiecomittee123):

In the following triangle, find the values of the angles B and B1, which are the best approximations to the solution to this ambiguous case.

OpenStudy (cutiecomittee123):



OpenStudy (cutiecomittee123):

I am so confused on how to even start here, please give me some pointers?

OpenStudy (campbell_st):

you can use the law of sines to find the value of B1 then B will be the supplementary angle

OpenStudy (cutiecomittee123):

Yes I see that but I am slightly confused here is what I have tried sin45/8.2=B1/10 can i just multiply both sides by ten to get B1?

OpenStudy (campbell_st):

well almost to find B1 use \[\frac{\sin(45)}{8.2} = \frac{\sin(B1)}{10}\] Your calculation well see B1 as an angle larger than 45

OpenStudy (campbell_st):

so then \[\sin(B1) = \frac{10 \times \sin(45)}{8.2}\]

OpenStudy (campbell_st):

i hope that helps you

OpenStudy (campbell_st):

or you could say \[B1 = \sin^{-1}(\frac{10 \times \sin(45)}{8.2})\]

OpenStudy (cutiecomittee123):

okay and that will find both angles then.

OpenStudy (cutiecomittee123):

whhy did you take inverse sin?

OpenStudy (campbell_st):

yes, but looking at the diagram you drew, I would Say B1 would be larger than 90.... and B is less than 90. the calculation above will only give an angle less than 90.

OpenStudy (cutiecomittee123):

Hmm so the angles are different? the diagram I drew is not to scale

OpenStudy (campbell_st):

if B1 is less than 90... then the triangle that has B as an interior angle isn't a triangle as it's an isosceles triangle. and you can't have 2 angles larger than 90

OpenStudy (cutiecomittee123):

Hmm well it does say that this is an ambiguous case

OpenStudy (campbell_st):

so the calculation above will give B and B1 will be the supplemetary. Does my explanation make sense?